OpenStudy (anonymous):

Estimate the instantaneous rate of change of y with respect to x at x =9 for y=-square root X

jimthompson5910 (jim_thompson5910):

Hint: y = -sqrt(x) y = -x^(1/2) y ' = -(1/2)x^(-1/2) y ' = -1/(2*sqrt(x))

OpenStudy (anonymous):

were did the 1/2 at the from and back come from for y ' = -(1/2)x^(-1/2)

OpenStudy (anonymous):

front*

jimthompson5910 (jim_thompson5910):

in general, sqrt(x) = x^(1/2)

jimthompson5910 (jim_thompson5910):

when you derive x^n you get n*x^(n-1) this is by the power rule

OpenStudy (anonymous):

ummm what does derive mean? I didn't learn anything about derivatives yet

jimthompson5910 (jim_thompson5910):

so that's how I went from y = -x^(1/2) to y ' = -(1/2)x^(-1/2)

jimthompson5910 (jim_thompson5910):

oh you haven't?

jimthompson5910 (jim_thompson5910):

that's odd considering how they want you to find the instantaneous rate of change

OpenStudy (anonymous):

no , we just learned how to get average and instant rate of change so far.

OpenStudy (anonymous):

it says estimate, I tried using the difference quotient, didn't work, I tried picking values close to 9 so x=9.001 and finding it through delta y divided by delta x, it didn't work

jimthompson5910 (jim_thompson5910):

well what you could do is find the average rate of change from x = 9 to x = 10 then you can find the average rate of change from x = 9 to x = 9.5 then you can find the average rate of change from x = 9 to x = 9.1 then you can find the average rate of change from x = 9 to x = 9.01 Notice the pattern? I'm getting the other x value closer and closer to 9 and this will estimate the instantaneous rate of change at x = 9

OpenStudy (anonymous):

but when I picked x=9.001 it didn't work. For another simpler question were x=7 and y=2/x, I tried picking 7.001 for x , as well as 7 and finding the average rate of change, but I got the wrong answer the question says to use technology, like a graphing calc

jimthompson5910 (jim_thompson5910):

yeah i recommend a graphing calculator as well because you can type in the function and then go to the table to input any value you want

jimthompson5910 (jim_thompson5910):

so you would type -sqrt(x) into y1 then go to the table and type in 9, 9.1, 9.01, 9.001 etc into the table to get your outputs

OpenStudy (anonymous):

similar*

jimthompson5910 (jim_thompson5910):

which you would use in the average rate of change formula given below Average Rate of Change from x1 to x2 = (f(x2) - f(x1))/(x2 - x1)

OpenStudy (anonymous):

Ok. I will do it and see. x=9.001 x=9 (9.001, -3.00017) (9, -3) y/x = -3.00017+3/9.001-9 =-0.17 The answer at the back of the book is -1/6. I guess that could be because I rounded, Thanks for you help.

OpenStudy (anonymous):

could you help with another question?

jimthompson5910 (jim_thompson5910):

well you're estimating, so -0.17 or -0.16667 is very close to what the true value is (-1/6) If you used the derivative, you would have found -1/6 exactly on the first try without any guesses or estimations, but you don't know about that yet

jimthompson5910 (jim_thompson5910):

sure

OpenStudy (anonymous):

two points P(1,1) and Q(x, 2x-x^2) lin on the curve y= 2x-x^2. Find the simplified expression for the slope of the secant PQ

jimthompson5910 (jim_thompson5910):

Use the slope formula m = (y2 - y1)/(x2 - x1)

jimthompson5910 (jim_thompson5910):

In this case x1 = 1 x2 = x y1 = 1 y2 = 2x - x^2

OpenStudy (anonymous):

I did, but then I didn't know what to do after since I had, 2x-x^2-1/x-1

jimthompson5910 (jim_thompson5910):

now factor the numerator

OpenStudy (anonymous):

(x+1)(x+1)/x-1 is that right?

jimthompson5910 (jim_thompson5910):

2x-x^2-1 -x^2+2x-1 -(x^2-2x+1) -(x-1)^2

jimthompson5910 (jim_thompson5910):

so (2x-x^2-1)/(x-1) turns into ( -(x-1)^2 )/(x-1) and that simplifies into -(x-1) -x + 1

jimthompson5910 (jim_thompson5910):

so the slope of secant PQ is -x+1

OpenStudy (anonymous):

oh ok. I was just factoring it wrong. ughh, so silly. Thank you so so so so so much.

jimthompson5910 (jim_thompson5910):

you're welcome